OpenStudy (thats_kyy):

Mike and his friends bought cheese wafers for $2 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $25 to buy a total of 20 packets of wafers of the two varieties.

OpenStudy (thats_kyy):

Part A: Write a system of equations that can be solved to find the number of packets of cheese wafers and the number of packets of chocolate wafers that Mike and his friends bought at the carnival. Define the variables used in the equations. (5 points) Part B: How many packets of chocolate wafers and cheese wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer. (5 points)

OpenStudy (thats_kyy):

@satellite73

OpenStudy (thats_kyy):

@Loser66

OpenStudy (thats_kyy):

@xapproachesinfinity

OpenStudy (anonymous):

solve \[x+y=20\\ 2x+y=25\] for \(x\) and \(y\)

OpenStudy (thats_kyy):

x= 10 y = 5

OpenStudy (anonymous):

no that is not right

OpenStudy (anonymous):

because \(10+5=15\) not \(20\)

OpenStudy (thats_kyy):

oh

OpenStudy (anonymous):

you have lots of ways to solve one way it to subtract the first equation from the second, you get \(x=5\) right away

OpenStudy (thats_kyy):

so x = 5 and y = 15

OpenStudy (thats_kyy):

@satellite73

OpenStudy (anonymous):

yes

OpenStudy (thats_kyy):

Ok So am I Done

OpenStudy (thats_kyy):

with the question @satellite73

OpenStudy (thats_kyy):

@sammixboo Help Please

OpenStudy (thats_kyy):

i Just Need Understanding Im An Online Student @sammixboo

sammixboo (sammixboo):

So did you come up with a system on equations to start with?

OpenStudy (thats_kyy):

No what do you mean by that im a slow sorry so can you explain

sammixboo (sammixboo):

Two equations to help you solve your question

OpenStudy (thats_kyy):

um yea like this x+y = 20 2x + y = 25 @sammixboo

OpenStudy (thats_kyy):

@sageyy

OpenStudy (thats_kyy):

@jim_thompson5910

sammixboo (sammixboo):

Right!

OpenStudy (thats_kyy):

okay So Whats My Next Step @sammixboo

OpenStudy (thats_kyy):

i solved it and got x= 5 y= 15

OpenStudy (thats_kyy):

@dan815